A compound known to be a pain reliever had the empirical formula C9H8O4. When a mixture of 5.02 mg of the unknown and 50.37 mg of camphor was prepared, the melting point of a portion of this mixture was determined. The observed melting point of the mixture was 156 °C. What is the molecular mass of this substance?


1
Expert's answer
2020-10-22T08:39:15-0400

Melting temperature of pure camphor = 178 °C

The cryoscopic constant k = 37.7 °C*kg/mol

"\Delta" T = 178-156 = 22 °C

"\Delta" T = k * n / m(camphor)

n = 22 * (50.37/1000)/37.7 = 0.0294 mol

M(exp) = 5.02/0.0294 = 171 g/mol (by experiment, approximate value)

M must be a multiple of 180 (12*9+8+16*4)

It means

M = 180 g/mol